OpenStudy (anonymous):

I need help with this. Please, I'll give out medals. (sqrt2)-sqrt(-2)) (sqrt(8)+sqrt(-2))

OpenStudy (anonymous):

\[(\sqrt{2}-\sqrt{2}) (\sqrt{8}-\sqrt{-2})\]